How many milliliters of pure acid must be added to 100 milliliters of a 10% acid solution to obtain a 20% solution?
Answer by ptaylor(1894)

Let x=amount of pure acid needed
Now we know that the amount of pure acid added (x) plus the amount of pure acid in the 100 ml solution (0.10*100) has to equal the amount of pure acid in the final solution (0.20(100+x)). So our equation to solve is:
x+0.10*100=0.20(100+x) get rid of parens (distributive law) and simplify
x+10=20+0.20x subtract 0.20x and also 10 from each side
x+10-10-0.20x=20-10+0.20x-0.20x collect like terms
0.80x=10 divide each side by 0.80
x=12.5 ml-----------------amount of pure acid needed
CK
12.5+10=0.20(100+12.5)=0.20(112.5)
22.5=22.5
Hope this helps----ptaylor